OpenStudy (anonymous):

f(x)=x^2+1 h(x)=x-2 1. find function of (foh)(x) 2. state the domain using interval notation

OpenStudy (welshfella):

(f o h) x can be found by replacing the 'x' in x^2 + 1 by h(x) ( that is x - 2) and simplifying the result

OpenStudy (welshfella):

so we have (x - 2)^2 + 1 simplify

OpenStudy (anonymous):

so x^2-4x+5

OpenStudy (welshfella):

yep

OpenStudy (anonymous):

ok how do I state the domain for that ?

OpenStudy (welshfella):

thats also correct. All polynomials have this domain.

OpenStudy (anonymous):

thank you....I get confused when the o is between the f and h or sometimes g and h

OpenStudy (anonymous):

I have 1 more if you could help me

OpenStudy (anonymous):

find the function of (gof)(x) simplify and state the domain for f(x)= 3/x g(x) =1/x

OpenStudy (anonymous):

I can do it in another question if you would like so I can give another badge

OpenStudy (welshfella):

ok

OpenStudy (anonymous):

ok I am closing this one and submitting it in a new question.